# Bifurcations of Highly Inclined Near Halo Orbits using Moser Regularization

### v1

@@ 3.4. Differential Correction and Continuation
We iterate until convergence, where a vector $X_{i}$ is considered to be a periodic orbit if $|F(X_{i})|<10^{-10}$. To compute the Jacobian matrix $DF$, we evaluate the linearized flow, as well as symbolically compute the derivatives of the dependent variables $\xi_{0},\eta_{0}$ and $\eta_{3}$ with respect to the free variables. See Appendix C for the details of this derivation.
@@
For the continuation of periodic orbits, we apply pseudo-arclength continuation [kuznetsov_applied_bifurcation_theory_2023, Section 10.2.1] in the unregularized system, while in the Moser regularized system we perform natural parameter continuation with respect to the Hamiltonian energy.
3.5. Vertical Collision Orbits
@@
Figure 4. Northern vertical collision orbit in Hill’s problem, shown in configuration space (left), the $(q_{3},p_{3})$ plane (center), and Moser-regularized coordinates $(\xi_{0},\eta_{0})$ (right). Red marker denotes apoapsis (maximum $q_{3}$), and the cross indicates the periapsis corresponding to collision with the light primary.
@@
4. Numerical Results
4.1. Bifurcations of Vertical Collision Orbits and Halo Orbits in Hill’s Problem
@@
Figure 5. Bifurcation diagram of the vertical collision orbit (central line) and the $L_{1}$ and $L_{2}$ halo families (branches) in Hill’s problem. The horizontal axis corresponds to the Hamiltonian energy $\hat{H}_{0}=h$, and the vertical axis shows the $x$-coordinate at the apoapsis.
Figure 6. Periodic orbit families bifurcating from vertical collision orbits (gray) in Hill’s problem: (1) butterfly family from the first period-doubling bifurcation; (2) $L_{1}$ and $L_{2}$ halo families from the first pitchfork bifurcation; (3) $r_{x}$-symmetric branches from the second pitchfork bifurcation; (4) doubly symmetric family from the second period-doubling bifurcation.
@@
Figure 7. Continuation of the $L_{1}$ halo family in Hill’s problem, from its bifurcation from the planar Lyapunov orbit to its termination in the vertical collision orbit.
4.2. Perturbation of Vertical Collision Orbits to CR3BP
We study the continuation of the vertical collision orbits of Hill’s problem under perturbation to CR3BP, with respect to the parameter $\nu=\mu^{1/3}$. The discussion here generalizes the similar description given for the KS regularization in [gomez_llibre_martinez_simo_dynamics_vol1, Section 2.3.6].
@@ 4.2. Perturbation of Vertical Collision Orbits to CR3BP
Figure 8. Displacements in $\xi_{2}$ (left) and $\eta_{1}$ (right) coordinates of the vertical collision orbit of Hill’s problem at apoapsis ($q_{3}$ maximal) and periapsis ($q_{3}=0$) under perturbation to halo orbits in the CR3BP.
@@
  1. (1)
    The vertical collision orbit is degenerate at two energy levels, $h=h_{1}$ and $h=h_{2}$. As an $r_{y}$-symmetric orbit, the orbit at $h=h_{2}$ is non-degenerate. Hence, the vertical collision orbit continues for $h\neq h_{1}$ to an $r_{y}$-symmetric orbit after perturbation to the CR3BP by Theorem 2.7.
  2. (2)
@@
4.3. Near-Collision Behavior of Halo Orbits in the Saturn-Enceladus System
@@
Figure 9. Bifurcation diagram of the $L_{1}$ and $L_{2}$ halo families in the Saturn-Enceladus system. The horizontal axis corresponds to the Jacobi energy $H_{\mu}=c$, and the vertical axis shows the $x$-coordinate at the apoapsis.
@@
Figure 10. Continuation of the $L_{1}$ (top) and $L_{2}$ (bottom) halo families in the Saturn-Enceladus CR3BP.
4.4. Bifurcation Surface of Halo Families from $\mu=0$ to $\mu=0.5$
@@
Figure 11. Bifurcation surfaces of the $L_{1}$ and $L_{2}$ halo families for mass ratios $\mu\in[0,0.5]$ and energy range $h\in[-1.0,0.4]$. The horizontal axis shows the $x$-coordinate at the apoapsis. Sections of the surface for selected planet-moon systems, as well as the Hill’s problem ($\mu=0$) and the Copenhagen problem ($\mu=0.5$), are highlighted. Red dotted curves indicate degeneracies: two associated with fold bifurcations and one with a pitchfork bifurcation at higher energy.
4.4.1. Earth-Moon Case
@@
Figure 12. Bifurcation diagram of the $L_{1}$ and $L_{2}$ halo families in the Earth-Moon system. The horizontal axis corresponds to the Jacobi energy $H_{\mu}=c$, and the vertical axis shows the $x$-coordinate at the apoapsis.
4.4.2. Copenhagen Case $(\mu=0.5)$
@@
Figure 13. Bifurcation diagram of the $L_{1}$ and $L_{2}$ halo families in the Copenhagen problem ($\mu=0.5$). The horizontal axis corresponds to the Jacobi energy $H_{\mu}=c$, and the vertical axis shows the $x$-coordinate at the apoapsis.
@@
Figure 14. Continuation of $L_{1}$ (top) and $L_{2}$ (bottom) halo orbits in the Copenhagen problem ($\mu=0.5$).
4.5. W4/W5 families in Hill’s Problem and Earth-Moon System
@@
Figure 15. Bifurcation diagram of the vertical collision orbit (central line) and the W4 and W5 families of orbits (branches) in Hill’s problem. The horizontal axis corresponds to the Hamiltonian energy $\hat{H}_{0}=h$, and the vertical axis shows the $y$-coordinate at the apoapsis.
Figure 16. Continuation of the W4 family in Hill’s problem.
@@
Figure 17. Bifurcation diagram of the $L_{1}$ halo orbits (central line) and the W4 and W5 families of orbits (branches) in the Earth-Moon system. The horizontal axis corresponds to the Hamiltonian energy $H_{\mu}=c$, and the vertical axis shows the $y$-coordinate at the apoapsis.
Figure 18. Continuation of the W4 family (green) in the Earth-Moon system, from its bifurcation from the $L_{1}$ halo orbit (blue) to its termination in the $L_{4}$ vertical orbit at $L_{4}$ (gray).
4.6. Period Doubling Bifurcation and Butterfly Orbit Family
The butterfly orbit family is a well-known class of periodic orbits catalogued in the JPL Periodic Orbit Catalogue [vaquero_poincare_2018]. These orbits share several characteristics with halo orbits, including a vertical geometry, close proximity to the light primary, and extended coverage of the polar region [grebow_multibody_2008]. Recently, butterfly orbits in the Saturn-Enceladus system have been proposed as candidate science orbits for a mission to Enceladus due to their coverage of the south polar plume region, which is of high scientific interest because of its active geysers [boone_approach_2024]. In the literature, butterfly orbits have been primarily studied in the Earth-Moon system, where they arise through a period doubling bifurcation of the $L_{2}$ halo family [howell-campbell_three-dimensional_1999, whitley_earth-moon_2018].
By continuation in the Moser regularized rescaled Hamiltonian $\hat{H}_{\mu}$, we obtain the following charaterization of the butterfly orbits:
  1. (1)
@@ 4.6. Period Doubling Bifurcation and Butterfly Orbit Family
Figure 19. Bifurcation diagrams of the butterfly orbit family in the Hill’s problem (left) and the Saturn-Enceladus CR3BP (right). The horizontal axis represents the Hamiltonian energy, while the vertical axis shows the $x$-coordinate at one of the two intersections with the Poincaré section $y=0$ (specifically, the intersection with the smaller $x$ value).
@@
Figure 20. Period-doubling bifurcation of the butterfly orbit family from a collisional $L_{2}$ halo orbit in the Saturn-Enceladus CR3BP.
4.7. Tri-Fly Orbits from Period Tripling Bifurcations
@@
Figure 21. Four tri-fly orbits which originate from a period-tripling bifurcation of a collisional $L_{2}$ halo orbits in the Saturn-Enceladus system.
@@
Figure 22 shows the bifurcation structure of the tri-fly families in the Hill’s problem and the Saturn-Enceladus systems. In the Hill’s problem, all four tri-fly families emerge simultaneously from the period-tripling bifurcation of the vertical collision orbit. Once we perturb to the CR3BP, the “up” and “down” families do not originate directly from this bifurcation point, but instead from a separate pitchfork bifurcation along the “left” tri-fly branch. This results in a sequential rather than a simultaneous emergence of the families.
Figure 22. Bifurcation graphs of the four tri-fly orbits in the Hill’s problem (left) and the Saturn-Enceladus CR3BP (right). For the “left” and “right” orbits, we plot the coordinates of the point with $p_{3}=0$ on the Poincaré section $y=0$. For the “up” and “down” orbits in the Hill’s problem, we plot the coordinates of the point with $p_{3}=0$ on the section $x=0$, which corresponds to the fixed-point locus of the symmetry $r_{x}$. For these orbits in the Saturn-Enceladus system, we plot the coordinates of the intersection point with $p_{3}=0$ and $z>0$, with $y>0$ for the “up” orbit and $y<0$ for the “down” orbit.
@@ Acknowledgements
Chankyu Joung and Otto van Koert were supported by the National Research Foundation of Korea (NRF), grant number (MSIT) (RS-2023-NR076656), funded by the Korean government. Chankyu Joung was partially supported by BK21 SNU Mathematical Sciences Division. The authors thank Cengiz Aydin for many helpful discussions during the course of this project. Chankyu Joung and Dayung Koh are grateful to Damon Landau at the Jet Propulsion Laboratory for his kind support and for valuable conversations related to Enceladus orbits. Chankyu Joung also thanks Dongho Lee and Beomjun Sohn for their help and assistance throughout the preparation of this work.
@@ Appendix B Restricted phase space
Example B.4 (Mathematical Pendulum).

### v2

@@ 3.4 Differential Correction and Continuation
We iterate until convergence, where a vector $X_{i}$ is considered to be a periodic orbit if $|F(X_{i})|$ is sufficiently small: we chose the threshold to be less than $10^{-10}$ for most computations, and switched to $10^{-30}$ for high-precision computations, which were needed for the investigation of touch-and-go bifurcations in Saturn-Enceladus. To compute the Jacobian matrix $DF$, we evaluate the linearized flow, as well as symbolically compute the derivatives of the dependent variables $\xi_{0},\eta_{0}$ and $\eta_{3}$ with respect to the free variables. See Appendix C for the details of this derivation.
@@
For the numerical continuation of periodic orbits, we apply pseudo-arclength continuation in the unregularized system, and switch to natural parameter continuation with respect to the Hamiltonian energy in the Moser-regularized system as trajectories approach the collision singularity. Initial periodic orbits are obtained either from the JPL Three-Body Periodic Orbit Catalog [JPL_periodic_orbits, vaquero_poincare_2018] or from internal data computed using the cell-mapping method [koh_cell_mapping_2021], and are transferred between systems and mass ratios using Algorithm 1. For several families studied in Sections 4.5 and 4.7, branch switching at local bifurcation points is applied. We refer to [kuznetsov_applied_bifurcation_theory_2023, Section 10.2] for the formulation of the continuation schemes and the branch-switching method.
3.5 Vertical Collision Orbits
@@
Figure 5: Northern vertical collision orbit in Hill’s problem, shown in configuration space (left), the $(q_{3},p_{3})$ plane (center), and Moser-regularized coordinates $(\xi_{0},\eta_{0})$ (right). Red marker denotes apoapsis (maximum $q_{3}$), and the cross indicates the periapsis corresponding to collision with the light primary.
@@
4 Numerical Results
4.1 Bifurcations of Vertical Collision Orbits and Halo Orbits in Hill’s Problem
@@
Figure 6: Bifurcation diagram of the vertical collision orbit (central line) and the $L_{1}$ and $L_{2}$ halo families (branches) in Hill’s problem. The horizontal axis corresponds to the Hamiltonian energy $\hat{H}_{0}=h$, and the vertical axis shows the $x$-coordinate at the apoapsis.
Figure 7: Periodic orbit families bifurcating from vertical collision orbits (gray) in Hill’s problem: (1) butterfly family from the first period-doubling bifurcation; (2) $L_{1}$ and $L_{2}$ halo families from the first pitchfork bifurcation; (3) $r_{x}$-symmetric branches from the second pitchfork bifurcation; (4) doubly symmetric family from the second period-doubling bifurcation.
@@
Figure 8: Continuation of the $L_{1}$ halo family in Hill’s problem, from its bifurcation from the planar Lyapunov orbit to its termination in the vertical collision orbit.
4.2 Perturbation of Vertical Collision Orbits to CR3BP
We study the continuation of the vertical collision orbits of Hill’s problem under perturbation to CR3BP, with respect to the parameter $\nu=\mu^{1/3}$. The discussion here extends the similar description given for the KS regularization in [gomez_llibre_martinez_simo_dynamics_vol1, Section 2.3.6].
@@ 4.2 Perturbation of Vertical Collision Orbits to CR3BP
Figure 9: Displacements in $\xi_{2}$ (left) and $\eta_{1}$ (right) coordinates of the vertical collision orbit of Hill’s problem at apoapsis ($q_{3}$ maximal) and periapsis ($q_{3}=0$) under perturbation to halo orbits in the CR3BP.
@@
  1. 1.
    The vertical collision orbit is degenerate at two energy levels, $h=h_{1}$ and $h=h_{2}$. As an $r_{y}$-symmetric orbit, the orbit at $h=h_{2}$ is non-degenerate. Hence, the vertical collision orbit continues for $h\neq h_{1}$ to an $r_{y}$-symmetric orbit after perturbation to the CR3BP by Theorem 4.
  2. 2.
@@
4.3 Near-Collision Behavior of Halo Orbits in the Saturn-Enceladus System
@@
Figure 10: Bifurcation diagram of the $L_{1}$ and $L_{2}$ halo families in the Saturn-Enceladus system. The horizontal axis corresponds to the Jacobi energy $H_{\mu}=c$, and the vertical axis shows the $x$-coordinate at the apoapsis.
@@
Figure 11: Continuation of the $L_{1}$ (top) and $L_{2}$ (bottom) halo families in the Saturn-Enceladus CR3BP.
4.4 Bifurcation Surface of Halo Families from $\mu=0$ to $\mu=0.5$
@@
Figure 12: Bifurcation surfaces of the $L_{1}$ and $L_{2}$ halo families for mass ratios $\mu\in[0,0.5]$ and energy range $h\in[-1.0,0.4]$. The horizontal axis shows the $x$-coordinate at the apoapsis. Sections of the surface for selected planet-moon systems, as well as the Hill’s problem ($\mu=0$) and the Copenhagen problem ($\mu=0.5$), are highlighted. Red dotted curves indicate degeneracies: two associated with fold bifurcations and one with a pitchfork bifurcation at higher energy.
4.4.1 Earth-Moon Case
@@
Figure 13: Bifurcation diagram of the $L_{1}$ and $L_{2}$ halo families in the Earth-Moon system. The horizontal axis corresponds to the Jacobi energy $H_{\mu}=c$, and the vertical axis shows the $x$-coordinate at the apoapsis.
4.4.2 Copenhagen Case $(\mu=0.5)$
@@
Figure 14: Bifurcation diagram of the $L_{1}$ and $L_{2}$ halo families in the Copenhagen problem ($\mu=0.5$). The horizontal axis corresponds to the Jacobi energy $H_{\mu}=c$, and the vertical axis shows the $x$-coordinate at the apoapsis.
@@
Figure 15: Continuation of $L_{1}$ (top) and $L_{2}$ (bottom) halo orbits in the Copenhagen problem ($\mu=0.5$).
4.5 W4/W5 families in Hill’s Problem and Earth-Moon System
@@
Figure 16: Bifurcation diagram of the vertical collision orbit (central line) and the W4 and W5 families of orbits (branches) in Hill’s problem. The horizontal axis corresponds to the Hamiltonian energy $\hat{H}_{0}=h$, and the vertical axis shows the $y$-coordinate at the apoapsis.
Figure 17: Continuation of the W4 family in Hill’s problem.
@@
Figure 18: Bifurcation diagram of the $L_{1}$ halo orbits (central line) and the W4 and W5 families of orbits (branches) in the Earth-Moon system. The horizontal axis corresponds to the Hamiltonian energy $H_{\mu}=c$, and the vertical axis shows the $y$-coordinate at the apoapsis.
Figure 19: Continuation of the W4 family (green) in the Earth-Moon system, from its bifurcation from the $L_{1}$ halo orbit (blue) to its termination in the $L_{4}$ vertical orbit at $L_{4}$ (gray).
4.6 Period Doubling Bifurcation and Butterfly Orbit Family
The butterfly orbit family is a well-known class of periodic orbits cataloged in the JPL Three-Body Periodic Orbit Catalog [JPL_periodic_orbits, vaquero_poincare_2018]. These orbits share several characteristics with halo orbits, including a vertical geometry, close proximity to the light primary, and extended coverage of the polar region [grebow_multibody_2008]. Recently, butterfly orbits in the Saturn-Enceladus system have been proposed as candidate science orbits for a mission to Enceladus due to their coverage of the south polar plume region, which is of high scientific interest because of its active geysers [boone_approach_2024]. In the literature, butterfly orbits have been primarily studied in the Earth-Moon system, where they arise through a period doubling bifurcation of the $L_{2}$ halo family [howell-campbell_three-dimensional_1999, whitley_earth-moon_2018].
By continuation in the Moser regularized rescaled Hamiltonian $\hat{H}_{\mu}$, we obtain the following characterization of the butterfly orbits:
  1. 1.
@@ 4.6 Period Doubling Bifurcation and Butterfly Orbit Family
Figure 20: Bifurcation diagrams of the butterfly orbit family in the Hill’s problem (left) and the Saturn-Enceladus CR3BP (right). The horizontal axis represents the Hamiltonian energy, while the vertical axis shows the $x$-coordinate at one of the two intersections with the Poincaré section $y=0$ (specifically, the intersection with the smaller $x$ value).
@@
Figure 21: Period-doubling bifurcation of the butterfly orbit family from a collisional $L_{2}$ halo orbit in the Saturn-Enceladus CR3BP.
4.7 Tri-Fly Orbits from Period Tripling Bifurcations
@@
Figure 22: Representative tri-fly orbits from the four branches near the triple cover of the $L_{2}$ halo orbit in the Saturn-Enceladus system.
@@
Figure 23 shows the bifurcation structure of the tri-fly families in Hill’s problem, the Saturn-Enceladus system, and the Earth-Moon system. In Hill’s problem, all four tri-fly families emerge simultaneously from the triple cover of the vertical collision orbit, with the four branches exhibiting a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry [aydin_Batkhin_hill_2025]. After perturbation to the CR3BP, this symmetry is broken. In the Saturn-Enceladus system, the “up” and “down” families bifurcate via a pitchfork bifurcation from the “left” tri-fly branch, whereas in the Earth-Moon system they bifurcate from the “right” branch. As a result, the four tri-fly families emerge sequentially rather than simultaneously.
Figure 23: Bifurcation graphs of the four tri-fly orbits in the Hill’s problem (left), Saturn-Enceladus system (center), and Earth-Moon system (right). Each graph tracks coordinates of a selected intersection point on the Poincaré section $p_{3}=0$. For the Saturn-Enceladus and Earth-Moon systems, the “left” and “right” branches follows the intersection with $y=0$, while for the “up” and “down” branches, the same point is continued as it bifurcates out of the plane $y=0$.
Figure 24 shows a zoomed view near the period-tripling bifurcation of the $L_{2}$ halo family for the Saturn-Enceladus and Earth-Moon systems. In both cases, two branches intersect transversely in the direction of increasing Jacobi energy, with both branches remaining on both sides of the energy at the bifurcation point. This is characteristic of a touch-and-go bifurcation [sadovskii_bifurcation_1996], also referred to as a phantom kiss [abraham_marsden_foundations_1978] in Meyer’s classification of generic bifurcations [meyer_generic_bifurcation_1970]. Immediately after the intersection, the “left” tri-fly branch undergoes a fold bifurcation, after which the Jacobi energy decreases. These observations are consistent with the computer-assisted proofs of touch-and-go bifurcations by Walawska and Wilczak [walawska_touch_and_go_2019] for $L_{2}$ halo orbits in the Earth-Moon and Sun-Jupiter systems. Such bifurcations occur on very small energy scales (on the order of $10^{-12}$ in the Saturn-Enceladus system). To reliably identify the bifurcation type at this scale, we employed for the Saturn-Enceladus system arbitrary-precision arithmetic [fousse_mpfr_2007] together with a high-order Taylor integrator implemented in the CAPD library [kapela_capd_2021]. The result confirms the presence of the same touch-and-go bifurcation in the Saturn-Enceladus system.
Figure 24: Zoomed-in bifurcation graphs near the period-tripling bifurcation of the $L_{2}$ halo family in the Saturn-Enceladus (left) and Earth-Moon (right) systems. The intersection of two branches in the direction of increasing Jacobi energy indicates a touch-and-go bifurcation.
Finally, we note that the “right” tri-fly orbit corresponds to the three-petal tulip-shaped orbit studied in [koblick_kelly_tulip_2025] for the Earth-Moon system. The remaining tri-fly families exhibit similar three-lobed spatial structures, differing primarily by orientation, and the corresponding orbits in Saturn-Enceladus involve close flybys of the south pole region of Enceladus (see Figure 25), making them potentially relevant for future mission design. From a dynamical viewpoint, the bifurcation mechanisms identified here provide insight into the organization of these orbit families. In the Earth-Moon system, we additionally observe a short-lived “spin” tri-fly branch, shown in brown in Figure 23, which connects the “left” and “right” branches over a narrow parameter range. Representative orbits from the “left”, “down”, and “spin” branches in the Earth-Moon system are shown in Figure 26.
Figure 25: Representative tri-fly orbits in the Saturn-Enceladus system, shown in nondimensional coordinates. The four trajectories belong to distinct branches located near the triple cover of the $L_{2}$ halo family. The orbits illustrated have minimum altitudes of approximately 30km above the surface of Enceladus (radius 252.1 km, shown in light blue).
Figure 26: Representative tri-fly orbits in the Earth-Moon system from the “left” (top), “down” (middle), and “spin” (bottom) branches. The black circle marks the location of the secondary primary. For the spin branch, the red cross indicates the initial point used for continuation.
@@ Acknowledgements
Chankyu Joung and Otto van Koert were supported by the National Research Foundation of Korea (NRF), grant number (MSIT) (RS-2023-NR076656), funded by the Korean government. Chankyu Joung was partially supported by BK21 SNU Mathematical Sciences Division. The authors thank Cengiz Aydin for many helpful discussions during the course of this project. Chankyu Joung and Dayung Koh are grateful to Damon Landau at the Jet Propulsion Laboratory for his kind support and for valuable conversations related to Enceladus orbits. Chankyu Joung also thanks researchers at the Jet Propulsion Laboratory for valuable interactions during his visit, as well as Dongho Lee and Beomjun Sohn for their assistance in the preparation of this work.
@@ Appendix B Restricted phase space
Example 1 (Mathematical Pendulum).